# High-dimensional covariance estimation by pairwise likelihood truncation

### v1

@@ 4.1 Monte Carlo experiments
In this section, we study the finite-sample properties of the TPL estimator focusing on its support-recovery performance. We draw $n$ i.i.d. samples from the $p$-variate normal distribution $N_{p}(\bf 0,{\bm{\Theta}})$ under different structures for ${\bm{\Theta}}$ by varying the sparsity level $\tau=(m{-}p{-}m_{0})/{{p}\choose{2}}$, i.e. the proportion of zero off-diagonal elements in ${\bm{\Theta}}$. The considered covariance structures are described as follows, and are also depicted in Figure 1. (i) Block diagonal structure: ${\bm{\Theta}}_{jj}=1$, for $j=1,\dots,p$, while off-diagonal elements are zero except for a block of adjacent variables, and whose dimension depends on $\tau$. The entries of the off-diagonal elements of the nonzero block are drawn independently from the univariate normal distribution $N(0.5,0.05^{2})$. Additionally, we also consider a first-order autoregressive model to determine the elements in the non-null block; the results are similar to those for the block diagonal structure and are, therefore, omitted here. (ii) Sparse at random structure. The non-zero entries of ${\bm{\Theta}}$ are generated at random from an Erdős-Rényi (ER) graph model with probability of connection $\tau$. The covariance matrix ${\bm{\Theta}}$ is obtained by finding a positive definite matrix closest to the sample covariance on the given ER random graph by the iterative conditional fitting algorithm of Chaduri et al. Chaudhuri et al. (2007). This structure is challenging for support recovery, since ${\bm{\Theta}}$ tends to have many nearly zero entries among the non-zero entries.
Figure 1: Examples of sparse at random (left) and block diagonal (right) structures for ${\bm{\Theta}}$ for $p=50$ and sparsity level $\tau=(m-p-m_{0})/{{p}\choose{2}}=0.9$. Black squares represent non-zero entries.
We consider $n=40,100,250$, $p=20,50,150$ and $\tau=0.5,0.9$. which allow us to explore both $n>p$ and $n<p$ scenarios. The performance of the method is explored in both sparse and relatively dense scenarios, with the latter somewhat deviating somewhat from the typical high-dimensional setting. For each combination, of $n$, $p$ and $\tau$, we generate $B=100$ Monte Carlo samples. The tuning constant $\lambda$ is selected according to (26) with $\alpha=0.1$ in all our simulations. The reported analyses have been conducted in the R environment (R Core Team, 2023), with some subroutines written in C++ to reduce the computing time. The support recovery performance of the TPL method is assessed using the following sensitivity (SN), specificity (SP) and accuracy (AC) metrics:
@@
Table 1 shows Monte Carlo estimates of SN, SP and AC statistics for both block diagonal and sparse at random structures. For the block diagonal structure, the support recovery looks generally favourable: all the metrics appear to be rather insensitive to changes in $n,p$ and $\tau$, attaining excellent results across all the settings. Particularly, the overall accuracy remains stable regardless of $\tau$ and $p$. While the specificity generally improves with $p$, a slight degradation in sensitivity is observed when $n=40$ and $p=150$. Interestingly, in this setting the TPLE estimator maintains desirable properties even with low sparsity ($\tau=0.5$), regardless of the data dimension, pointing out the applicability of the methodology in scenarios beyond sparse estimation.
Interesting observations can be drawn from the more challenging sparse at random setting, where many elements nearly zero make it harder to reconstruct the exact covariance structure. Although specificity generally increases with $p$, sensitivity is not entirely satisfactory, especially with $\tau=0.5$ and $p$ large. While possibly disappointing, this behaviour is linked to the relative efficiency of the estimator. Particularly, scenarios, where the sensitivity is lower, correspond also to ones where $\hat{{\bm{\Theta}}}$ outperforms the oracle ML estimator ${\bm{\Theta}}^{oml}$, thus indicating a bias-variance trade-off. In this setting, with the current naive choice of $\lambda$ (depending on a fixed $\chi^{2}(1)$ quantile), our method tends to over-sparsify, especially in dense scenarios, by shrinking small but non-zero elements to zero. This results in many false negatives, reducing sensitivity. However, over-sparsification significantly reduces variance, thus reducing the overall mean squared error.
@@ 4.2 Mouse brain single-cell RNA data
Single-cell transcriptome sequencing (scRNA-seq) is a powerful technique that enables RNA sequencing from individual cells and allows the study of their specific gene expression profiles (Kolodziejczyk et al., 2015). We illustrate the TPL estimation procedure by analyzing a dataset collected by Zeisel et al. Zeisel et al. (2015) to provide a census of two brain regions, namely the hippocampal CA1 and the primary somatosensory cortex. The dataset is available in the Gene Expression Omnibus database (accession number GSE60361) and consists of $n=2816$ mouse brain cells for which the expression of $p=19839$ genes has been measured. In our analysis, we focus on two types of immune cells, microglia and perivascular macrophages, corresponding to $n_{\texttt{mgl}}=28$ and $n_{\texttt{pvm}}=50$ observations, respectively. Whilst originating from from distinct developmental pathways, these two cell types are closely related and play a crucial role in brain development, homeostasis regulation and neurodegenerative diseases. In this context, sparse covariance estimates can disentangle complex gene interactions and offer visual summaries through their link to covariance graph models and gene co-expression networks (Chaudhuri et al., 2007). Data quality checks are conducted following Amezquita et al. Amezquita et al. (2020), while data normalization, pre-processing and feature filtering steps follow Townes et al. Townes et al. (2019). The number of considered genes is therefore reduced to $p=200$. According to the literature, this is a conservative overestimate, as the number of informative genes in scRNA-seq is typically relatively small within a single tissue, due to limited biological variation.
@@
In Figure 2, we show covariance estimates for the two types of cells obtained by the TPL estimator. Consistently with our experiments on simulated data, we use $\alpha=0.1$ to select $\lambda$ according to the criterion in (26). A first visual inspection highlights how the estimated sparsity structures are different. Although we considered the same $\alpha$, the sparsity pattern looks strongly dependent on the cell type. Particularly, the number of non-zero off-diagonal elements corresponds to 7.89% and 13.93% for microglia and perivascular macrophages types, respectively. The non-zero correlations are generally positive, while negative correlations are, in most cases, shrunk to zero. The estimated matrices differ not only in terms of sparsity magnitude, but also in terms of the location of the non-zero elements. In Figure 3, we show estimated covariance matrices for microglia for $\alpha=0.01$ and $\alpha=0.4$, which correspond to 2.58% and 26.98% non-zero elements in $\hat{\bm{\Theta}}$, respectively. This illustrates how the proposed thresholding approach can be used to explore different requirements in terms of sparsity magnitude by tuning $\alpha$. In turn, this is a starting point for further explorations which can provide some insight on the relationships among the genes in the considered cells.
@@ 5 Conclusion and final remarks
Despite recent methodological advancements in high-dimensional covariance estimation, significant challenges remain in constructing unbiased estimators that achieve optimal statistical accuracy while ensuring model selection consistency when $p$ is much larger than $n$. The primary contribution of this work is the development of a framework for high-dimensional covariance selection and estimation, integrating some distinctive traits of penalized likelihood and thresholding methodologies. The selection rule introduced in Equations (7)–(9) penalizes the discrepancy between the PL and ML score functions, thus aiming to maximize the statistical accuracy encoded by the PL score function for a given level of sparsity. Similar to Lasso-type approaches in regression, our method leverages the geometric properties of the $L_{1}$ penalty to set certain terms of the PL function to zero (see Theorem 1). The new criterion also functions as a form of adaptive thresholding, adjusting the marginal information that determines selection uncertainty based on the chosen covariance estimates (see Section 2.3).
@@
A key distinction of our TPL procedure from existing penalization approaches for high-dimensional estimation is its focus on selecting terms in the pairwise likelihood function, with each term corresponding to an entire bivariate model, rather than penalizing individual elements of ${\bm{\Theta}}$ in a likelihood criterion. While penalized likelihood estimators rely on biased estimating equations to induce selection Fan and Li (2001); Zou (2006), an advantage of our construction is that the resulting pairwise estimating equations are inherently unbiased. Therefore, conditional on the consistent selection of the relevant variable pairs provided in Theorem 4, the resulting covariance estimates remain unbiased, and the estimator $\hat{\bm{\Theta}}$ behaves asymptotically as the oracle ML estimator (Corollary 1). In terms of model selection performance, we showed that the TLP procedure has the selection consistency property when $p$ grows exponentially in $n$, and the data generating process has arbitrary exponential tails. The selection consistency result stated in Theorem 4 is generally stronger than sparsistency alone - i.e., the zero covariances are correctly estimated with probability tending to one – with the latter implied by Part (i) of Theorem 4. Since the conditions outlined in Theorem 4 also ensure power converging to one, they are more stringent than necessary to prove sparsistency alone.
@@
The form of the solution given in Theorem 1 and the discussion in Sections 2.3 reveals a connection between PL estimation and adaptive thresholding frameworks that, to our knowledge, has not been previously explored. Specifically, the criterion function in Equation (7) introduces a general method for constructing adaptive thresholding procedures, depending on the form of the pairwise score and the form of the penalty. For instance, when the variance elements on the diagonal of ${\bm{\Theta}}$ are fixed, the score covariance matrix $\hat{\mathbf{J}}$ becomes diagonal, meaning that the pairwise scores contain information on separate covariance parameters. Consequently, the resulting thresholding rule matches the adaptive thresholding method of Cai and Liu Cai and Liu (2011). In future research, sparse covariance estimation in the context of heavy-tailed data distributions may be achieved using pairwise scores based on bivariate t-distributions. Similarly, estimating the dependence structure for high-dimensional extreme events could be explored using bivariate extreme value distributions.
@@
The numerical results presented in Section 4 corroborate the support recovery properties of the TPL procedure for sparse matrices in finite samples for different covariance structures. Additionally, they provide insight into the behaviour of the estimator in the challenging case of sparse at random, unstructured covariance with relatively low sparsity. The presence of many nearly zero covariance elements leads to general overshrinkage based on our naive choice of $\lambda$ based on the chi-square 0.9-quantile. While this reduces the sensitivity of the method, it also helps avoid the accumulation of estimation errors, resulting in mean squared error considerably smaller than that of the oracle ML estimator, suggesting that further improvement is possible by some form of adaptive selection of $\lambda$. In this regard, the selection criterion in Equation (26) suggests treating the selection of $\lambda$ as a top-down sequential hypothesis testing procedure for nested models. Following a logic similar to that of Lockhart et al. Lockhart et al. (2014) in the context of sparse linear regression, one can order hypotheses according to a set of knots for $\lambda$, such as $\lambda_{max}>\dots>\lambda_{min}\geq 0$, each corresponding to the entry of a new pairwise score. Starting from $\lambda_{max}$ corresponding to a diagonal $\hat{\mathbf{\Theta}}$, pairwise scores are added at each knot until a relatively dense estimate $\hat{\mathbf{\Theta}}$ is reached at $\lambda_{min}$. This opens avenues for future research on the interplay between hyperparameter selection and sequential model selection procedures, particularly in controlling false discovery rates in ordered testing settings G’Sell et al. (2016).
@@
Note that ${\mathbf{u}}_{jk}({\bm{\theta}};{\bm{X}})$ is the only sub-likelihood score with the element in postion $\{(2p+2-j)(j-1)/2+k+1-j\}$ containing $\theta_{jk}$ for $j<k$. Since ${\bm{\Theta}}$ is positive definite, we have $\theta_{jk}^{2}<\theta_{jj}\theta_{kk}$. This implies that ${\mathbf{u}}_{jk}$ is a nonzero vector with probability one. By Formula (3) and (4), all bivariate scores ${\mathbf{u}}_{jk}$ are linearly independent. Thus, by the rank equality of Gram matrices, $\hat{\mathbf{J}}$ and $\hat{\mathbf{J}}_{\hat{\varepsilon}}$ are both full ranked. This implies that $\hat{\mathbf{w}}$ exists and is unique due to the strict convexity of the first term of Criterion $\hat{d}_{\lambda}({\mathbf{w}})$ (13) and the convexity of the second (penalty) term. Let ${\mathbf{u}}_{jk}^{(i)}={\mathbf{u}}_{jk}({\mathbf{s}};{\bm{X}}^{(i)})$ and ${\mathbf{u}}^{(i)}={\mathbf{u}}({\mathbf{s}},\hat{\mathbf{w}};{\bm{X}}^{(i)})$. By the Karush-Kuhn-Tucker (KKT) conditions for quadratic optimization, the solution must satisfy
@@
Under Conditions C1-C4, for any constant $r_{1}\in(0,\frac{r}{8+r})$ and $r_{2}=\min(\frac{r}{32+4r},\frac{4+r}{32+4r}-\frac{r_{1}}{4})$, let $p=o\{\exp(n^{r_{1}-c})\}$ and $m_{0}=o(n^{r_{2}-c})$ for some arbitrarily small $c\in(0,\min\{r_{1},r_{2}\})$, then $\|{\mathbf{S}}-{\bm{\Theta}}\|_{max}=o_{p}(m_{0}^{-2})$.
@@
Under Conditions C1-C4, for any constant $r_{1}\in(0,\frac{r}{8+r})$ and $r_{2}=\min(\frac{r}{32+4r},\frac{4+r}{32+4r}-\frac{r_{1}}{4})$, let $m_{0}=o(n^{r_{2}-c})$ for some arbitrarily small $c\in(0,\min\{r_{1},r_{2}\})$, then we have $\|\hat{\mathbf{J}}_{\xi}-{\mathbf{J}}_{\xi}\|_{max}=o_{p}(m_{0}^{-2})$.
@@
Under Conditions C1-C4, for any constant $r_{1}\in(0,\frac{r}{8+r})$ and $r_{2}=\min(\frac{r}{32+4r},\frac{4+r}{32+4r}-\frac{r_{1}}{4})$, let $m_{0}=o(n^{r_{2}-c})$ for some arbitrarily small $c\in(0,\min\{r_{1},r_{2}\})$. If $\Pr(\hat{\mathcal{E}}\subseteq{\mathcal{E}})\to 1$ as $n\to\infty$, then $\|\hat{\mathbf{w}}_{\xi}\|_{1}=O_{p}(m_{0})$ and $\hat{\mathbf{w}}_{{\mathcal{E}}\setminus\xi}$ is a vector of ones with probability converging to 1.
@@
From the objective (13), if the lemma assumption holds, then $\Pr(\|\hat{\mathbf{w}}\|_{1}\leq\|\tilde{\mathbf{w}}_{\mathcal{E}}\|_{1})\to 1$ as $n\to\infty$, where $\tilde{\mathbf{w}}_{\mathcal{E}}={\hat{\mathbf{J}}_{{\mathcal{E}}}}^{-1}\hat{\mathbf{h}}_{{\mathcal{E}}}$ is a length $|{\mathcal{E}}|$ vector optimizing ${\mathbf{w}}^{\top}\hat{\mathbf{J}}_{\mathcal{E}}{\mathbf{w}}/2-{\mathbf{w}}^{\top}\hat{\mathbf{h}}_{\mathcal{E}}$. Let ${\mathbf{w}}_{\mathcal{E}}^{\ast}={\mathbf{J}}_{\mathcal{E}}^{-1}{\mathbf{h}}_{\mathcal{E}}$ be the minimizer of ${\mathbf{w}}^{\top}{\mathbf{J}}_{\mathcal{E}}{\mathbf{w}}/2-{\mathbf{w}}^{\top}{\mathbf{h}}_{\mathcal{E}}$. Then $\hat{\mathbf{J}}_{\mathcal{E}}\tilde{\mathbf{w}}_{\mathcal{E}}=\hat{\mathbf{h}}_{\mathcal{E}}$ and ${\mathbf{J}}_{\mathcal{E}}{\mathbf{w}}_{\mathcal{E}}^{\ast}={\mathbf{h}}_{\mathcal{E}}$. These two equations imply that ${\mathbf{J}}_{\mathcal{E}}(\tilde{\mathbf{w}}_{\mathcal{E}}-{\mathbf{w}}_{\mathcal{E}}^{\ast})=\hat{\mathbf{h}}_{\mathcal{E}}-{\mathbf{h}}_{\mathcal{E}}+({\mathbf{J}}_{\mathcal{E}}-\hat{\mathbf{J}}_{\mathcal{E}})\tilde{\mathbf{w}}_{\mathcal{E}}$. Moreover, by the definition of 14, $\hat{\mathbf{J}}_{\mathcal{E}}{\mathbf{w}}=((\hat{\mathbf{J}}_{\xi}{\mathbf{w}}_{\xi})^{\top},(\hat{\mathbf{J}}_{{\mathcal{E}}\setminus\xi}{\mathbf{w}}_{{\mathcal{E}}\setminus\xi})^{\top})^{\top}$ for any length $|{\mathcal{E}}|$ vector ${\mathbf{w}}$. I.e., the equation can be separate into two, each corresponding to the subset $\xi$ and ${\mathcal{E}}\setminus\xi$. The same statement holds for ${\mathbf{J}}_{\mathcal{E}}$. By the KKT Condition 28, this implies that $\hat{\mathbf{w}}_{{\mathcal{E}}\setminus\xi}=\tilde{\mathbf{w}}_{{\mathcal{E}}\setminus\xi}={\mathbf{w}}_{{\mathcal{E}}\setminus\xi}^{\ast}$ are vectors of ones, since $\hat{\mathbf{J}}_{{\mathcal{E}}\setminus\xi}$ and ${\mathbf{J}}_{{\mathcal{E}}\setminus\xi}$ are diagonal by definition when $\hat{\mathcal{E}}\subseteq{\mathcal{E}}$. Thus $\Pr(\|\hat{\mathbf{w}}_{\xi}\|_{1}\leq\|\tilde{\mathbf{w}}_{\xi}\|_{1})\to 1$ as $n\to\infty$.

### v2

@@ 4.1 Monte Carlo experiments
In this section, we study the finite-sample properties of the TPL estimator focusing on its support-recovery performance. We draw $n$ i.i.d. samples from the $p$-variate normal distribution $N_{p}(\bf 0,{\bm{\Theta}})$ under different structures for ${\bm{\Theta}}$ by varying the sparsity level $\tau$, defined as the proportion of zero off-diagonal elements in ${\bm{\Theta}}$. The considered covariance structures are described as follows, and are also depicted in Figure 1. (i) Block diagonal structure: ${\bm{\Theta}}_{jj}=1$, for $j=1,\dots,p$, while off-diagonal elements are zero except for a block of adjacent variables, and whose dimension depends on $\tau$. The entries of the off-diagonal elements of the nonzero block are drawn independently from the univariate normal distribution $N(0.5,0.05^{2})$. Additionally, we also consider first-order autoregressive model to determine the elements in the non-null block; the results are similar to those for the block diagional structure and are therefore omitted here. (ii) Sparse at random structure. The non-zero entries of ${\bm{\Theta}}$ are generated at random from an Erdős-Rényi (ER) graph model with probability of connection $\tau$. The covariance matrix ${\bm{\Theta}}$ is obtained by finding a positive definite matrix closest to the sample covariance on the given ER random graph by the iterative conditional fitting algorithm of Chaduri et al. Chaudhuri et al. (2007). This structure is challenging for support recovery, since ${\bm{\Theta}}$ tends to have many nearly zero entries among the non-zero entries.
Figure 1: Examples of sparse at random (left) and block diagonal (right) structures for ${\bm{\Theta}}$ for $p=50$ and sparsity level $\tau=0.9$. Black squares represent non-zero entries.
We consider $n=40,100,250$, $p=20,50,150$ and $\tau=0.5,0.9$. which allow us to explore both $n>p$ and $n<p$ scenarios. The performance of the method explored in both sparse and relatively dense scenarios, with the latter deviating somewhat from the typical high-dimensional setting. For each combination of $n$, $p$ and $\tau$, we generate $B=100$ Monte Carlo samples. The tuning constant $\lambda$ is selected according to (26) with $\alpha=0.1$ in all our simulations. The reported analyses have been conducted in the R environment (R Core Team, 2023), with some subroutines written in C++ to reduce the computing time. The support recovery performance of the TPL method is assessed using the following sensitivity (SN), specificity (SP) and accuracy (AC) metrics:
@@
Table 1 shows Monte Carlo estimates of SN, SP and AC statistics for both block diagonal and sparse at random structures. For the block diagonal structure, the support recovery looks generally favorable: all the metrics appear to be rather insensitive to changes in $n,p$ and $\tau$, attaining excellent results across all the settings. Particularly, the overall accuracy remains stable regardless of $\tau$ and $p$. While the specificity generally improves with $p$, a slight degradation in sensitivity is observed when $n=40$ and $p=150$. Interestingly, in this setting the TPL estimator maintains desirable properties even with low sparsity ($\tau=0.5$), regardless of the data dimension, pointing out the applicability of the methodology in scenarios beyond sparse estimation.
Interesting observations can be drawn from the more challenging sparse at random setting, where many elements nearly zero make it harder to reconstruct the exact covariance structure. Although specificity generally increases with $p$, sensitivity is not entirely satisfactory, especially with $\tau=0.5$ and $p$ large. While possibly disappointing, this behavior is linked to the relative efficiency of the estimator. Particularly, scenarios where the sensitivity is lower correspond also to ones where $\hat{{\bm{\Theta}}}$ outperforms the oracle ML estimator $\hat{\bm{\Theta}}^{oml}$, our method tends to over-sparsify, especially in dense scenarios, by shrinking small but non-zero elements to zero. This results in many false negatives, reducing sensitivity. However, over-sparsification significantly reduces variance, thus reducing the overall mean squared error.
@@ 4.2 Mouse brain single-cell RNA data
Single-cell transcriptome sequencing (scRNA-seq) is a powerful technique that enables RNA sequencing from individual cells and allows the study of their specific gene expression profiles (Kolodziejczyk et al., 2015). We illustrate the TPL estimtion procedure by analyzing a dataset collected by Zeisel et al. Zeisel et al. (2015) to provide a census of two brain regions, namely the hippocampal CA1 and the primary somatosensory cortex. The dataset is available in the Gene Expression Omnibus database (accession number GSE60361) and consists of $n=2816$ mouse brain cells for which the expression of $p=19839$ genes has been measured. In our analysis, we focus on two types of immune cells, microglia and perivascular macrophages, corresponding to $n_{\texttt{mgl}}=28$ and $n_{\texttt{pvm}}=50$ observations, respectively. Although they originate from distinct developmental pathways, these two cell types are closely related and play a crucial role in brain development, homeostatsis regulation and neurodegenerative diseases. In this context, sparse covariance estimates can disentangle complex gene interactions and offer visual summaries through their link to covariance graph models and gene co-expression networks (Chaudhuri et al., 2007). Data quality checks are conducted following Amezquita et al. Amezquita et al. (2020), while data normalization, pre-processing and feature filtering steps follow Townes et al. Townes et al. (2019). The number of considered genes is therefore reduced to $p=200$. According to the literature, this is a conservative overestimate, as the number of informative genes in scRNA-seq is typically relatively small within a single tissue, due to limited biological variation.
@@
In Figure 2, we show covariance estimates for the two types of cells obtained by the TPL estimator. Consistently with our experiments on simulated data, we use $\alpha=0.1$ to select $\lambda$ according to the criterion in (26). A first visual inspection highlights how the estimated sparsity structures are different. Although we considered the same $\alpha$, the sparsity pattern looks strongly dependent on the cell type. Particularly, the number of non-zero off-diagonal elements corresponds to 7.89% and 13.93% for microglia and perivascular macrophages types, respectively. The non-zero correlations are generally positive, while negative correlations are, in most of the cases, shrunk to zero. The estimated matrices differ not only in terms of sparsity magnitude, but also in terms of the location of the non-zero elements. In Figure 3, we show estimated covariance matrices for microglia for $\alpha=0.01$ and $\alpha=0.4$, which correspond to 2.58% and 26.98% non-zero elements in $\hat{\bm{\Theta}}$, respectively. This illustrates how the proposed thresholding approach can be used to explore different requirements in terms of sparsity magnitude by tuning $\alpha$. In turn, this is a starting point for further explorations which can provide some insight on the relationships among the genes in the considered cells.
@@ 5 Conclusion and final remarks
Despite recent methodological advancements in high-dimensional covariance estimation, significant challenges remain in constructing unbiased estimators achieving optimal statistical accuracy, while ensuring model selection consistency when $p$ is much larger than $n$. The primary contribution of this work is the development of a framework for high-dimensional covariance selection and estimation, featuring some distinctive traits of penalized likelihood and thresholding methodologies. The selection rule introduced in Equations (7)–(9) penalizes the discrepancy between the PL and ML score functions, thus aiming to maximize the statistical accuracy encoded by the PL score function for a given level of sparsity. Similar to Lasso-type approaches in regression, our method leverages the geometric properties of the $L_{1}$ penalty to set certain terms of the PL function to zero (see Theorem 1). The new criterion also functions as a form of adaptive thresholding, adjusting the marginal information that determines selection uncertainty based on the chosen covariance estimates (see Section 2.3).
@@
A key distinction of our TPL procedure from existing penalization approaches for high-dimensional estimation is its focus on selecting terms in the pairwise likelihood function, with each term corresponding to an entire bivariate model, rather than penalizing individual elements of ${\bm{\Theta}}$ in a likelihood criterion. While penalized likelihood estimators rely on biased estimating equations to induce selection Fan and Li (2001); Zou (2006), an advantage of our construction is that the resulting pairwise estimating equations are inherently unbiased. Therefore, conditional on consistent selection of the relevant variable pairs provided in Theorem 4, the resulting covariance estimates remain unbiased and the estimator $\hat{\bm{\Theta}}$ behaves asymptotically as the oracle ML estimator (Corollary 1). In terms of model selection performance, we showed that the TLP procedure has the selection consistency property when $p$ grows exponentially in $n$ and the data generating process has arbitrary exponential tails. The selection consistency result stated in Theorem 4 is generally stronger than sparsistency alone - i.e., the zero covariances are correctly estimated with probability tending to one – with the latter implied by Part (i) of Theorem 4. Since the conditions outlined in Theorem 4 also ensure power converging to one, they are more stringent than necessary to prove sparsistency alone.
@@
The form of the solution given in Theorem 1 and the discussion in Sections 2.3 reveals a connection between PL estimation and adaptive thresholding frameworks that, to our knowledge, has not been previously explored. Specifically, the criterion function in Equation (7) introduces a general method for constructing adaptive thresholding procedures, depending on the distributional assumption for the pairwise score and the type of the penalty. For instance, when the variance elements on the diagonal of ${\bm{\Theta}}$ are fixed, the score covariance matrix $\hat{\mathbf{J}}$ becomes diagonal, meaning that the pairwise scores contain information on separate covariance parameters. Consequently, the resulting thresholding rule matches the adaptive thresholding method of Cai and Liu Cai and Liu (2011). In future research, sparse covariance estimation in the context of heavy-tailed data distributions may be achieved using pairwise scores based on bivariate t-distributions. Similarly, estimating the dependence structure for high-dimensional extreme events could be explored using bivariate extreme value distributions.
@@
The numerical results presented in Section 4 corroborate the support recovery properties of the TPL procedure for sparse matrices in finite samples for different covariance structures. Additionally, they provide insight into the behavior of the estimator in the challenging case of sparse at random unstructured covariance with relatively low sparsity. The presence of many nearly zero covariance elements leads to general overshrinkage based on our naive choice of $\lambda$ based on the chi-square 0.9-quantile. While reducing the sensitivity of the method, this also helps avoid the accumulation of estimation errors, resulting in mean squared error considerably smaller than that of the oracle ML estimator. Overall this behavior suggests that further improvements are possible through adaptive selection of $\lambda$. Particulary, the selection criterion in Equation (26) suggests treating the selection of $\lambda$ as a top-down sequential hypothesis testing procedure for nested models. Following the approach of Lockhart et al. Lockhart et al. (2014) in the context of sparse linear regression, one can order hypotheses according to a set of knots for $\lambda$, such as $\lambda_{max}>\dots>\lambda_{min}\geq 0$, each corresponding to the entry of a new pairwise score. Starting from $\lambda_{max}$ corresponding to a diagonal $\hat{\mathbf{\Theta}}$, pairwise scores are added at each knot until a relatively dense estimate $\hat{\mathbf{\Theta}}$ is reached at $\lambda_{min}$. This opens avenues for future research on the interplay between hyperparameter selection and sequential model selection procedures, particularly in controlling false discovery rates in ordered testing settings G’Sell et al. (2016).
@@
Note that ${\mathbf{u}}_{jk}({\bm{\theta}};{\bm{X}})$ is the only sub-likelihood score with the element in position $\{(2p+2-j)(j-1)/2+k+1-j\}$ containing $\theta_{jk}$ for $j<k$. Since ${\bm{\Theta}}$ is positive definite, we have $\theta_{jk}^{2}<\theta_{jj}\theta_{kk}$. This implies that ${\mathbf{u}}_{jk}$ is a nonzero vector with probability one. By Formula (3) and (4), all bivariate scores ${\mathbf{u}}_{jk}$ are linearly independent. Thus, by the rank equality of Gram matrices, $\hat{\mathbf{J}}$ and $\hat{\mathbf{J}}_{\hat{\varepsilon}}$ are both full ranked. This implies that $\hat{\mathbf{w}}$ exists and is unique due to the strict convexity of the first term of Criterion $\hat{d}_{\lambda}({\mathbf{w}})$ (13) and the convexity of the second (penalty) term. Let ${\mathbf{u}}_{jk}^{(i)}={\mathbf{u}}_{jk}({\mathbf{s}};{\bm{X}}^{(i)})$ and ${\mathbf{u}}^{(i)}={\mathbf{u}}({\mathbf{s}},\hat{\mathbf{w}};{\bm{X}}^{(i)})$. By the Karush-Kuhn-Tucker (KKT) conditions for quadratic optimization, the solution must satisfy
@@
Under Conditions C1-C4, for any constant $r_{1}\in(0,\frac{r}{8+r})$ and $r_{2}=\min(\frac{r}{32+4r},\\ \frac{4+r}{32+4r}-\frac{r_{1}}{4})$, let $p=o\{\exp(n^{r_{1}-c})\}$ and $m_{0}=o(n^{r_{2}-c})$ for some arbitrarily small $c\in(0,\min\{r_{1},r_{2}\})$, then $\|{\mathbf{S}}-{\bm{\Theta}}\|_{max}=o_{p}(m_{0}^{-2})$.
@@
Under Conditions C1-C4, for any constant $r_{1}\in(0,\frac{r}{8+r})$ and $r_{2}=\min(\frac{r}{32+4r},\\ \frac{4+r}{32+4r}-\frac{r_{1}}{4})$, let $m_{0}=o(n^{r_{2}-c})$ for some arbitrarily small $c\in(0,\min\{r_{1},r_{2}\})$, then we have $\|\hat{\mathbf{J}}_{\xi}-{\mathbf{J}}_{\xi}\|_{max}=o_{p}(m_{0}^{-2})$.
@@
Under Conditions C1-C4, for any constant $r_{1}\in(0,\frac{r}{8+r})$ and $r_{2}=\min(\frac{r}{32+4r},\\ \frac{4+r}{32+4r}-\frac{r_{1}}{4})$, let $m_{0}=o(n^{r_{2}-c})$ for some arbitrarily small $c\in(0,\min\{r_{1},r_{2}\})$. If $\Pr(\hat{\mathcal{E}}\subseteq{\mathcal{E}})\to 1$ as $n\to\infty$, then $\|\hat{\mathbf{w}}_{\xi}\|_{1}=O_{p}(m_{0})$ and $\hat{\mathbf{w}}_{{\mathcal{E}}\setminus\xi}$ is a vector of ones with probability converging to 1.
@@
From the objective (13), if the lemma assumption holds, then $\Pr(\|\hat{\mathbf{w}}\|_{1}\leq\|\tilde{\mathbf{w}}_{\mathcal{E}}\|_{1})\to 1$ as $n\to\infty$, where $\tilde{\mathbf{w}}_{\mathcal{E}}={\hat{\mathbf{J}}_{{\mathcal{E}}}}^{-1}\hat{\mathbf{h}}_{{\mathcal{E}}}$ is a length $|{\mathcal{E}}|$ vector optimizing ${\mathbf{w}}^{\top}\hat{\mathbf{J}}_{\mathcal{E}}{\mathbf{w}}/2-{\mathbf{w}}^{\top}\hat{\mathbf{h}}_{\mathcal{E}}$. Let ${\mathbf{w}}_{\mathcal{E}}^{\ast}={\mathbf{J}}_{\mathcal{E}}^{-1}{\mathbf{h}}_{\mathcal{E}}$ be the minimizer of ${\mathbf{w}}^{\top}{\mathbf{J}}_{\mathcal{E}}{\mathbf{w}}/2-{\mathbf{w}}^{\top}{\mathbf{h}}_{\mathcal{E}}$. Then $\hat{\mathbf{J}}_{\mathcal{E}}\tilde{\mathbf{w}}_{\mathcal{E}}=\hat{\mathbf{h}}_{\mathcal{E}}$ and ${\mathbf{J}}_{\mathcal{E}}{\mathbf{w}}_{\mathcal{E}}^{\ast}={\mathbf{h}}_{\mathcal{E}}$. These two equations imply that ${\mathbf{J}}_{\mathcal{E}}(\tilde{\mathbf{w}}_{\mathcal{E}}-{\mathbf{w}}_{\mathcal{E}}^{\ast})=\hat{\mathbf{h}}_{\mathcal{E}}-{\mathbf{h}}_{\mathcal{E}}+({\mathbf{J}}_{\mathcal{E}}-\hat{\mathbf{J}}_{\mathcal{E}})\tilde{\mathbf{w}}_{\mathcal{E}}$. Moreover, by the definition of 14, $\hat{\mathbf{J}}_{\mathcal{E}}{\mathbf{w}}=((\hat{\mathbf{J}}_{\xi}{\mathbf{w}}_{\xi})^{\top},(\hat{\mathbf{J}}_{{\mathcal{E}}\setminus\xi}{\mathbf{w}}_{{\mathcal{E}}\setminus\xi})^{\top})^{\top}$ for any length $|{\mathcal{E}}|$ vector ${\mathbf{w}}$. I.e., the equation can be separated into two, each corresponding to the subset $\xi$ and ${\mathcal{E}}\setminus\xi$. The same statement holds for ${\mathbf{J}}_{\mathcal{E}}$. By the KKT Condition 28, this implies that $\hat{\mathbf{w}}_{{\mathcal{E}}\setminus\xi}=\tilde{\mathbf{w}}_{{\mathcal{E}}\setminus\xi}={\mathbf{w}}_{{\mathcal{E}}\setminus\xi}^{\ast}$ are vectors of ones, since $\hat{\mathbf{J}}_{{\mathcal{E}}\setminus\xi}$ and ${\mathbf{J}}_{{\mathcal{E}}\setminus\xi}$ are diagonal by definition when $\hat{\mathcal{E}}\subseteq{\mathcal{E}}$. Thus $\Pr(\|\hat{\mathbf{w}}_{\xi}\|_{1}\leq\|\tilde{\mathbf{w}}_{\xi}\|_{1})\to 1$ as $n\to\infty$.